# From small eigenvalues to large cuts, and Chowla’s cosine problem

### v1

@@ 13 Edit distance from the union of cliques
The limitations of Theorems 9.4 and 9.6 are that they are only meaningful if the graph has large density. Indeed, these theorems imply that a graph of large minimum eigenvalue (or small surplus) is $o(1)$-close to a disjoint union of cliques, where a careful inspection of the parameters show that the $o(1)$ term is polylogarithmic in $n$. Hence, if the graph $G$ has density at most $n^{-c}$ for any small constant $c>0$, they become meaningless as $G$ is already $o(1)$-close to the empty graph. However, in this section we prove that polynomial proximity can also be establish, albeit under somewhat stronger condition on the smallest eigenvalue or the surplus.
We start with the following simple lemma, which will be used to argue that a dense graph with small surplus cannot induce sparse subgraphs.
@@ 13.1 Least eigenvalue version
Let $s\in(0,1/6)$ and $\delta>2s$, then there exists $\varepsilon>0$ such that the following holds for sufficiently large $n$. Let $G$ be a graph on $n$ vertices such that $|\lambda_{n}|\leq n^{s}$. Then there exists $X\subset V(G)$ such that the number of edges not in $G[X]$ is at most $n^{2-\varepsilon}$, and $G[X]$ can be partitioned into cliques of size $n^{1-\delta}$.
@@
Choose any constants $s_{0},\delta_{0}>0$ such that $s<s_{0}$ and $2s_{0}+\delta_{0}<\delta$. By Theorem 12.1, there exists $\rho=\rho(s_{0},\delta_{0})>0$ such that every $n_{0}$-vertex graph of edge density at least $n_{0}^{-\rho}$ and smallest eigenvalue at least $n_{0}^{s_{0}}$ contains a clique of size $n_{0}^{1-2s_{0}-\delta_{0}}$ as long as $n_{0}$ is sufficiently large.
We show that $\varepsilon=\min\{1/16,\rho/5,(1-s/s_{0})/2,(\delta-\delta_{0}-2s_{0})/2\}$ suffices. Delete all vertices of $G$ of degree less than $d=n^{1-2\varepsilon}$, and let $G_{0}$ be the resulting graph. Note that we removed at most $dn$ edges. Repeat the following procedure. If $G_{i}$ is defined and $G_{i}$ contains a clique of size $n^{1-\delta}$, then let $C_{i+1}$ be such a clique, and set $G_{i+1}=G_{i}\setminus C_{i+1}$. Otherwise, stop, and let $I$ be the last index $i$ for which $G_{i}$ is defined. We show that $X=C_{1}\cup\dots\cup C_{I}$ is the required set. It is clear that $X$ can be partitioned into cliques of size $n^{1-\delta}$, so it remains to show that the number of edges not in $X$ is at most $n^{2-\varepsilon}$.
Let $Y=V(G_{I})$, $b=e(G_{0}[X,Y])$ and $c=e(G_{0}[Y])$.
@@
As $G$ has minimum degree $d$, we have $b+2c\geq d|Y|$. Assume that $c<d^{2}|Y|^{2}/(20n^{2})<d|Y|/4$, then $b\geq d|Y|/2$. By Lemma 13.1, we have
$\operatorname{surp}^{*}(G)\geq\operatorname{surp}^{*}(G_{0})\geq\frac{b^{2}}{4n^{2}}-c.$
Using that $\operatorname{surp}^{*}(G)\leq|\lambda_{n}|n\leq n^{1+s}$, we have
$c\geq\frac{b^{2}}{4n^{2}}-n^{1+s}\geq\frac{d^{2}|Y|^{2}}{16n^{2}}-n^{1+s}\geq\frac{d^{2}|Y|^{2}}{20n^{2}}.$
In the last inequality we used that $d=n^{1-2\varepsilon}\geq n^{7/8}$ and $|Y|\geq n^{3/4}$. ∎
Now we show that $|Y|\leq n^{1-2\varepsilon}$. Indeed, suppose $|Y|>n^{1-2\varepsilon}$. By the above claim, $G_{I}$ has at least $d^{2}|Y|^{2}/(20n^{2})>|Y|^{2}n^{-4\varepsilon}/20\geq|Y|^{2-5\varepsilon}$ edges, which shows that $G_{I}$ has edge density at least $|Y|^{-5\varepsilon}\geq|Y|^{-\rho}$. Moreover, by Cauchy’s interlacing theorem, the smallest eigenvalue of $G_{I}$ is at least that of $G$, which is at least $n^{s}\geq-|Y|^{s/(1-2\varepsilon)}\geq-|Y|^{s_{0}}$.
But then, as discussed above, Theorem˜12.1 guarantees that $Y$ contains a clique of size $|Y|^{1-2s_{0}-\delta_{0}}>n^{1-2s_{0}-\delta_{0}-2\varepsilon}\geq n^{1-\delta}$, contradicting that $G_{I}$ contains no clique of size $n^{1-\delta}$. Therefore, we must have $|Y|\leq n^{1-2\varepsilon}$.
From this, the number of edges of $G$ not in $G[X]$ is at most
$dn+e(G[X,Y])+e(G[Y])\leq dn+|Y|n\leq n^{1-\varepsilon}.$
This finishes the proof. ∎
Next, we show that the graph between two cliques must be either very dense or very sparse. In order to prove this, we use that graphs of small $|\lambda_{n}|$ avoid the following simple graph as an induced subgraph.
@@
By the Cauchy interlacing theorem and Claim 3.3, $G$ does not contain $H_{k}$ as an induced subgraph for $k=2|\lambda_{n}|^{2}$. We may assume that $|X|=|Y|\geq 4k$, otherwise the statement is trivial. Then each vertex in $X$ has either at most $k$ neighbours or at most $k$ non-neighbours in $Y$. Moreover, each vertex in $Y$ has at most $k$ neighbours or at most $k$ non-neighbours in $X$. Let $X_{0}\subset X$ be the set of vertices with at most $k$ neighbours, and let $X_{1}=X\setminus X_{0}$, and define $Y_{0},Y_{1}\subset Y$ analogously. Suppose that $X_{0}$ and $Y_{1}$ both have size at least $2c|\lambda_{n}|^{2}$. If the number of edges between $X_{0}$ and $Y_{1}$ is at least $|X_{0}||Y_{1}|/2$, then there is a vertex in $X_{0}$ with at least $|Y_{1}|/2\geq k$ neighbours in $Y_{1}$, contradiction. On the other hand, if the number of edges between $X_{0}$ and $Y_{1}$ is at most $|X_{0}||Y_{1}|/2$, then there is a vertex in $Y_{0}$ with at least $|X_{0}|/2\geq k$ non-neighbours, contradiction. Therefore, it must hold that at least one of $X_{0}$ or $Y_{1}$ has size at most $2k$. If $|X_{0}|\leq 2k$, then $G[X,Y]$ has at least $|X_{1}|(|X|-k)\geq|X|^{2}-3k|X|=|X|^{2}-O(|\lambda_{n}|^{2}|X|)$ edges. Otherwise, if $|Y_{1}|\leq 2k$, then $G[X,Y]$ has at most $|Y_{0}|k+|Y_{1}||X|\leq 3k|X|=O(|\lambda_{n}|^{2}|X|)$ edges. ∎
@@
Let $s\in(0,1/6)$, then there exists $\varepsilon>0$ such that the following holds. Let $G$ be an $n$-vertex graph such that $|\lambda_{n}|\leq n^{s}$. Then $G$ is $n^{-\varepsilon}$-close to the disjoint union of cliques.
@@
Let $\delta=1/2>2s$, and let $\varepsilon_{0}=\varepsilon_{0}(s,\delta)>0$ be the constant guaranteed by Lemma 13.2. We show that $\varepsilon=\min\{1/7,\varepsilon_{0}/2\}$ suffices.
By Lemma 13.2, there exists a set $X\subset V(G)$ such that $X$ can be partitioned into the union of cliques of size $n_{0}=n^{1-\delta}$, and $G$ has at most $n^{2-\varepsilon_{0}}$ edges not in $G[X]$. Let $C_{1},\dots,C_{I}$ be the cliques of size $n_{0}$ partitioning $X$, then $I=|X|/n_{0}$. Lemma 13.4 implies that the bipartite graph between $C_{i}$ and $C_{j}$ has either at most $O(n_{0}|\lambda_{n}|^{2})$, or at least $n_{0}^{2}-O(n_{0}|\lambda_{n}|^{2})$ edges. In other words, $(C_{i},C_{j})$ is a $O(|\lambda_{n}|^{2}/n_{0})$-homogeneous pair, whose definition appears in Section˜9. Here, $O(|\lambda_{n}|^{2}/n_{0})<n^{-1/6}$. Define the auxiliary graph $\Gamma$ on vertex set $\{1,\dots,I\}$, where we connect $i$ and $j$ if $(C_{i},C_{j})$ is $n^{-1/6}$-full.
@@
$\Gamma$ contains no cherry.
@@
If $\Gamma$ contains a cherry, it means that there is a triple $(C_{i},C_{j},C_{k})$ such that $(C_{i},C_{j})$ and $(C_{j},C_{k})$ are $n^{-1/6}$-full, but $(C_{i},C_{k})$ is $n^{-1/6}$-empty. By Lemma 9.2, then $G[C_{i}\cup C_{j}\cup C_{k}]$ has surplus at least $(1/4-3n^{-1/6})|C_{i}|^{2}\geq|C_{i}|n^{1/2}/8$, which shows that the smallest eigenvalue of $G[C_{i}\cup C_{j}\cup C_{k}]$ is at most $n^{1/2}/24$, contradiction. ∎
@@
Recall that graphs containing no cherry are the disjoint union of cliques. Therefore, we can partition $V(\Gamma)$ into sets $I_{1},\dots,I_{s}$ such that $\Gamma[I_{i}]$ is a clique and there are no edges between $I_{i}$ and $I_{j}$ in $\Gamma$. But this gives a partition of $X$ into sets $Y_{1},\dots,Y_{s}$ by setting $Y_{a}=\bigcup_{i\in I_{a}}C_{i}$. Define $\widetilde{G}$ to be the graph on vertex set $V(G)$, where $Y_{1},\dots,Y_{s}$ are cliques, and all edges of $\widetilde{G}$ are contained in one of these cliques.
@@
We prove that $\widetilde{G}$ is $n^{-\varepsilon}$-close to $G$. For $1\leq i<j\leq I$, $G[C_{i},C_{j}]$ and $\widetilde{G}[C_{i},C_{j}]$ differ by at most $n_{0}^{2}\cdot n^{-1/6}$ edges. Therefore, $G[X]$ and $\widetilde{G}[X]$ differ by at most
$\binom{|X|/n_{0}}{2}n_{0}^{2}\cdot n^{-1/6}\leq n^{2-1/6}$

### v2

@@ 11 Edit distance from the union of cliques
Theorems˜1.4 and 8.7 become meaningless once the graph $G$ has density at most $n^{-c}$ for any small constant $c>0$, as then $G$ is already $o(1)$-close to the empty graph. However, in this section we prove Theorem˜1.5, which deals with these graphs, under a somewhat stronger condition on the smallest eigenvalue.
The proof works as follows. Using Theorem˜10.1, we can repeatedly pull out “large” cliques in $G$ as long as the rest graph has “many” edges. Then, we show that the union of these cliques induce almost all edges in $G$, so it is sufficient to consider this disjoint union of cliques. Moreover, we show that between any two cliques, it is either very sparse or very dense. We use this to derive that $G$ itself must resemble a disjoint union of cliques.
We start with the following simple lemma, which will be used to argue that a dense graph with large smallest eigenvalue cannot induce sparse subgraphs ($G[Y]$ in the following statement). In fact, we can prove a stronger statement in terms of the surplus (recall that $\operatorname{surp}(G)\leq n|\lambda_{n}|/4$ from ˜5.1).
@@
Let $\gamma\in(0,1/6)$, then there exists $\alpha>0$ such that the following holds for sufficiently large $n$. Let $G$ be a graph on $n$ vertices such that $|\lambda_{n}|\leq n^{\gamma}$. Then there exists $X\subset V(G)$ such that the number of edges not in $G[X]$ is at most $n^{2-\alpha}$, and $G[X]$ can be partitioned into cliques of size $\sqrt{n}$.
@@
Choose any constants $\gamma_{0},\delta_{0}>0$ such that $\gamma_{0}>\gamma$, $2\gamma_{0}+\delta_{0}<3/8$ and $6\gamma_{0}+\delta_{0}<1$. Then, there exists a constant $\rho\in(0,1/2)$ such that $\rho/\delta_{0}+2\gamma_{0}/(1-\delta_{0}-4\gamma_{0})<1$. By Theorem 10.1, every $n_{0}$-vertex graph of edge density at least $n_{0}^{-\rho}$ and smallest eigenvalue at least $n_{0}^{\gamma_{0}}$ contains a clique of size $n_{0}^{1-2\gamma_{0}-\delta_{0}}\geq n_{0}^{5/8}$ as long as $n_{0}$ is sufficiently large.
We show that $\alpha=\min\{1/16,\rho/5,(1-\gamma/\gamma_{0})/2\}$ suffices. Repeatedly delete vertices of degree less than $d=n^{1-2\alpha}$, and let $G_{0}$ be the resulting graph. Note that we removed at most $dn$ edges. We are done if $G_{0}$ is empty, so we may assume $G_{0}$ has minimum degree at least $d$. Let $X$ be the maximal subset of $V(G)$ which can be partitioned into disjoint cliques of size $\sqrt{n}$.
[…]
$dn+e(G[X,Y])+e(G[Y])\leq dn+|Y|n\leq n^{1-\alpha}.$
Suppose, for the sake of contradiction, that $|Y|>n^{1-2\alpha}$. By maximality of $X$, the set $Y$ contains no clique of size $\sqrt{n}$, and we will ultimately derive the contradiction to this.
@@
Following the notation in the Lemma˜11.1, let $b=e(G_{0}[X,Y])$ and $c=e(G_{0}[Y])$. As $G_{0}$ has minimum degree $d$, we have $b+2c\geq d|Y|$. If $c\geq\frac{1}{4}d|Y|$, then the density of $G[Y]$ is at least $\frac{1}{4}d|Y|/\binom{|Y|}{2}\geq\frac{d}{2|Y|}\geq\frac{n^{1-2\alpha}}{2|Y|}>|Y|^{-5\alpha}\geq|Y|^{-\rho}$. Now, we assume that $c<\frac{1}{4}d|Y|$, so $b>\frac{1}{2}d|Y|$. By Lemma˜11.1, we have
$\operatorname{surp}(G)\geq\operatorname{surp}(G_{0})\geq\frac{b^{2}}{4n^{2}}-c.$
On the other hand, $\operatorname{surp}(G)\geq\frac{1}{4}|\lambda_{n}|n\leq n^{1+\gamma}$ by ˜5.1, so
$c\geq\frac{b^{2}}{4n^{2}}-n^{1+\gamma}>\frac{d^{2}|Y|^{2}}{16n^{2}}-n^{1+\gamma}\geq\frac{d^{2}|Y|^{2}}{20n^{2}}.$
In the last inequality we used that $d,|Y|\geq n^{1-2\alpha}$. Hence, $G[Y]$ has at least $e(G[Y])\geq d^{2}|Y|^{2}/(20n^{2})\geq|Y|^{2}n^{-4\alpha}/20\geq|Y|^{2-5\alpha}\geq|Y|^{2-\rho}$ edges, which shows that $G_{I}$ has edge density at least $|Y|^{-\rho}$. ∎
By Cauchy’s interlacing theorem, the smallest eigenvalue of $G[Y]$ is at least that of $G$, i.e. we have $|\lambda_{n}(G[Y])|\leq|\lambda_{n}(G)|\leq n^{\gamma}$. Since $|Y|>n^{1-2\alpha}$, this shows $|\lambda_{n}(G[Y])|\leq|Y|^{\gamma/(1-2\alpha)}\leq|Y|^{\gamma_{0}}$.
But then, as discussed in the beginning, Theorem˜10.1 guarantees that $Y$ contains a clique of size $|Y|^{5/8}>n^{(5/8)(1-2\alpha)}\geq\sqrt{n}$, contradicting the assumption that $Y$ contains no clique of size $\sqrt{n}$. Therefore, we must have $|Y|\leq n^{1-2\alpha}$, and as discussed above, this finishes the proof. ∎
Next, we show that the graph between two cliques must be either very dense or very sparse.
@@
Recall that in Section 3, we identified a graph $H_{k}$, consisting of a clique of size $2k$ and an additional vertex with $k$ neighbours in the clique, with the property that no graph $G$ with at $\lambda_{n}\geq-\sqrt{k/2}$ contains $H_{k}$ as an induced subgraph (see Claim 3.3 and the subsequent discussion). Thus, if we set $k=2|\lambda_{n}|^{2}$, the graph $G$ does not contain $H_{k}$ as an induced subgraph. We may assume that $|X|=|Y|\geq 4k$, otherwise the statement is trivial. Then each vertex in $X$ has either at most $k$ neighbours or at most $k$ non-neighbours in $Y$. Moreover, each vertex in $Y$ has at most $k$ neighbours or at most $k$ non-neighbours in $X$. Let $X_{0}\subset X$ be the set of vertices with at most $k$ neighbours, and let $X_{1}=X\setminus X_{0}$, and define $Y_{0},Y_{1}\subset Y$ analogously. Suppose that $X_{0}$ and $Y_{1}$ both have size at least $2k$. If the number of edges between $X_{0}$ and $Y_{1}$ is at least $|X_{0}||Y_{1}|/2$, then there is a vertex in $X_{0}$ with at least $|Y_{1}|/2\geq k$ neighbours in $Y_{1}$, contradiction. On the other hand, if the number of edges between $X_{0}$ and $Y_{1}$ is at most $|X_{0}||Y_{1}|/2$, then there is a vertex in $Y_{0}$ with at least $|X_{0}|/2\geq k$ non-neighbours, contradiction. Therefore, it must hold that at least one of $X_{0}$ or $Y_{1}$ has size at most $2k$. If $|X_{0}|\leq 2k$, then $G[X,Y]$ has at least $|X_{1}|(|X|-k)\geq|X|^{2}-3k|X|=|X|^{2}-O(|\lambda_{n}|^{2}|X|)$ edges. Otherwise, if $|Y_{1}|\leq 2k$, then $G[X,Y]$ has at most $|Y_{0}|k+|Y_{1}||X|\leq 3k|X|=O(|\lambda_{n}|^{2}|X|)$ edges. ∎
@@
For every $\gamma\in(0,1/6)$, there exists $\alpha>0$ such that for every sufficiently large $n$ we have the following. If $G$ is an $n$-vertex graph with $|\lambda_{n}|\leq n^{\gamma}$, then $G$ is $n^{-\alpha}$-close to the vertex-disjoint union of cliques.
@@
Let $\alpha_{0}=\alpha_{0}(\gamma)>0$ be the constant guaranteed by Lemma˜11.2. We show that $\alpha=\min\{1/7,\alpha_{0}/2\}$ works.
By Lemma˜11.2, there exists a set $X\subset V(G)$ such that $X$ can be partitioned into the union of cliques of size $\sqrt{n}$, and $G$ has at most $n^{2-\alpha_{0}}$ edges not in $G[X]$. Let $C_{1},\dots,C_{I}$ be the cliques of size $\sqrt{n}$ partitioning $X$; then $I=|X|/\sqrt{n}$. Lemma˜11.4 implies that the bipartite graph between $C_{i}$ and $C_{j}$ has either at most $O(\sqrt{n}|\lambda_{n}|^{2})$, or at least $n-O(\sqrt{n}|\lambda_{n}|^{2})$ edges. If $n$ is sufficiently large, $O(|\lambda_{n}|^{2}/\sqrt{n})<n^{-1/6}$. Define the auxiliary graph $\Gamma$ on vertex set $\{1,\dots,I\}$, where we connect $i$ and $j$ if $G[C_{i},C_{j}]$ has density at least $1-n^{-1/6}$.
@@
$\Gamma$ contains no cherry, i.e. if $ij,jk\in E(\Gamma)$, then $ik\in E(\Gamma)$ as well.
@@
If there is a triple $(C_{i},C_{j},C_{k})$ such that $G[C_{i},C_{j}],G[C_{j},C_{k}]$ have density at least $1-n^{-1/6}$, but $G[C_{i},C_{k}]$ has density at most $1-n^{-1/6}$, then we can apply Lemma˜8.3. This lemma shows that $\operatorname{surp}(G[C_{i}\cup C_{j}\cup C_{k}])\geq(1/4-3n^{-1/6})|C_{i}|^{2}\geq n/8$. Then, ˜5.1 implies the smallest eigenvalue of $G[C_{i}\cup C_{j}\cup C_{k}]$ is at most $n^{1/2}/6$, which is a contradiction since $|\lambda_{n}(G[C_{i}\cup C_{j}\cup C_{k}])|\leq|\lambda_{n}(G)|=n^{\gamma}$. ∎
@@
Recall that graphs containing no cherry are the disjoint union of cliques. Therefore, we can partition $V(\Gamma)$ into sets $I_{1},\dots,I_{\ell}$ such that $\Gamma[I_{a}]$ is a clique and there are no edges between $I_{a}$ and $I_{b}$ in $\Gamma$. But this gives a partition of $X$ into sets $Y_{1},\dots,Y_{\ell}$ by setting $Y_{a}=\bigcup_{i\in I_{a}}C_{i}$. Define $\widetilde{G}$ to be the graph on vertex set $V(G)$, where $Y_{1},\dots,Y_{\ell}$ are cliques, and all edges of $\widetilde{G}$ are contained in one of these cliques.
@@
We prove that $\widetilde{G}$ is $n^{-\alpha}$-close to $G$. For $1\leq i<j\leq I$, $G[C_{i},C_{j}]$ and $\widetilde{G}[C_{i},C_{j}]$ differ by at most $n\cdot n^{-1/6}=n^{5/6}$ edges. Therefore, $G[X]$ and $\widetilde{G}[X]$ differ by at most
$\binom{|X|/\sqrt{n}}{2}\cdot n^{5/6}\leq n^{11/6}$